Find the magnitude and direction angle θ of the following vectors. Round the magnitude to the nearest tenth, and round the direction angle to the nearest degree, if rounding is necessary.
v = (-5, -2)

Answer by Theo(13342)

if this vector starts at the origin, then you will get:

p1 = (0,0)
p2 = (-5,-2)

the x-coordinate of p2 is -5
the y-coordinate of p2 is -2

those coordinates form a right triangle with the origin.

the tangent of the angle of that triangle would be -2/-5.

tan^-1(2/5) = 21.80140949 degrees.
that's the reference angle of the triangle in quaedrant 3.
the angle in quadrant 3 is 180 + 21.80140949 = 201.80140949 degrees.

that's the direction.

the magnitude is sqrt((-2)^2 + (-5)^2) = sqrt(4+25) = sqrt(29).

the direction in quadrant 3 is generally in a south west direction.

since the reference angle is only 21 degrees, it's closer to west than south.

halfway south and halfway west would be an angle of 180 + 45 = 225 degrees.